OpenStudy (anonymous):

Renata is completing a craft project that involves covering only the lateral surface of a cylindrical container with fabric. The cylinder has a height of 12.8 in. and a radius of 15.1 in. To the nearest square unit, how much fabric does she need for this project? Use a calculator. A. 1,214 in.2 B. 2,040 in.2 C. 2,647 in.2 D. 1,931 in.2

zepdrix (zepdrix):

Hmm, so what type of problem is this? Volume? Surface Area? Can you tell based on the way it's worded? :)

OpenStudy (anonymous):

wellll is said surdace

zepdrix (zepdrix):

Yesss good :D Do you remember the formula for Surface Area of a Cylinder? \[\huge A=2\pi r^2+2\pi r h\] But in this problem, we only care about the LATERAL area, so we'll ignore the area of the top and bottom.\[\huge A=2\pi r h\]

zepdrix (zepdrix):

From here it shouldn't be too bad! Just plug in the numbers they gave you. :D Lemme know if you're stuck on any part.

OpenStudy (anonymous):

so is 2pi*r*h

zepdrix (zepdrix):

yes, that will give you the amount of lateral area that you need to cover with fabric. they gave you -r and h- in the question. If your calculator doesn't have a PI button, use 3.14, you should still get close to the correct answer.